A ball with projection velocity 10 metre per second has to hit a target 6 m away in the same horizontal line. how high above the target must the ball be aimed so that the ball will hit the target.

Grade:11

1 Answers

Pawan joshi
5 years ago
Height of the gun will be equal to vertical displacement of the ball.... .Also angle made by gun with horizontal will be so small such that we can neglect the initial vertical velocity, ie., Uy≈0.
time taken to cover distance =6/10s
Vertical displacement =Uyt+1/2at^2
y=0+1/2×10×(6/10)^2=1.8m